# EvalTree: Profiling Language Model Weaknesses via Hierarchical Capability Trees

### v1

@@ E.3.2 Analysis on Experimental Results
To better understand why TextDiff and QualEval are outperformed, we show the Precision and Recall curves in Figure 11 and 12. These curves show that both methods suffer from poor Precision, indicating that the weaknesses they identify cannot pinpoint where the LM fails precisely. We present the identified weaknesses from TextDiff, QualEval, and EvalTree when the weakness profile size is 10 in Table 28, along with their corresponding Precision, Recall, and F1; they are based on applying the three methods to the synthetic evaluation result generated for the MATH benchmark, with the probability hyperparameters set to $p=0.7$ and $d=0.2$. We observe that EvalTree achieves significantly higher Precision compared to the other two methods, while maintaining a quite high Recall, indicating that EvalTree can more precisely pinpoint specific areas where the LM underperforms and thus better satisfy desideratum 1. For example, EvalTree identified the weakness “Analyzing and applying relationships among polynomial expressions and their roots using Vieta’s formulas”, which closely aligns with the ground-truth weakness “Solving polynomial equations by analyzing relationships through Vieta’s formulas”; in contrast, TextDiff and QualEval identified two much coarser-grained weaknesses, “Handling problems involving the properties of polynomials and their roots” and “Solving linear, polynomial, and quadratic equations, including factoring and roots” respectively, failing to capture the critical aspect of Vieta’s formulas.
@@ E.3.3 Computing F1 on a Separate Set
In this subsection, we present the results of Section 5.2 using a separate test set (distinct from the profiling set) for computing $A$ in the formulas outlined in Appendix E.3.1.
@@
Figure 13: F1 score curves of TextDiff, QualEval, and EvalTree, with the weakness profile size varying from 1 to 20. Precision, Recall, and thus F1 (more specifically, $A$ in the formulas outlined in Appendix E.3.1) are computed on a separate test set, distinct from the profiling set used to generate the synthetic evaluation results. A horizontal line indicates each method’s highest score. $d$ is a hyperparameter to control the sampling probability (see Appendix E.3.1).
@@ E.4 Experimental Details of Extrinsic Assessment
We use OpenAI’s gpt-4o-mini-2024-07-18 (OpenAI, 2024a) in our experiments to generate data inputs; the input generation prompts for MATH (Hendrycks et al., 2021b) and DS-1000 (Lai et al., 2023) are provided in Table 29 and Table 30, respectively; we set the max new tokens and temperature to 4096 and 1.0 (for generation diversity), respectively. We also use gpt-4o-mini-2024-07-18 to generate outputs for each collected input; the output generation prompts for MATH and DS-1000 are provided in Table 31 and Table 32, respectively; we set the max new tokens and temperature to 4096 and 0.0, respectively.
@@
For the EvalTree-guided data collection strategy, we set the accuracy threshold $\tau$ to 0.4 in the node extraction algorithm described in Section 3.2. This resulted in 9 identified weaknesses for MATH and 5 for DS-1000; the same number of weaknesses was identified when using the TextDiff-guided strategy and the QualEval-guided strategy, ensuring that all weakness-guided data collection strategies use weakness profiles of the same size. When sampling five in-context examples for input generation given an identified weakness in a weakness-guided data collection strategy, the examples are sampled from the associated instances of the identified weakness in the TextDiff-guided strategy, from the instances assigned to the identified weakness in the QualEval-guided strategy, and from the instances linked to the corresponding node in the EvalTree-guided strategy.
For each data collection strategy, we collect 128 synthetic instances for training. We finetune the models using LoRA (Hu et al., 2022), with a rank of 256, an alpha of 512, and a dropout rate of 0.1. The batch size is fixed at 8, and the maximum sequence length is set to 1024 tokens. Training is conducted using BF16 precision. The optimizer is configured with a learning rate of 1E-4, a cosine learning rate scheduler, a warmup ratio of 0.1, and no weight decay. The models are trained for 3 and 2 epochs in the experiments on MATH and DS-1000, respectively. These configurations are applied consistently across all experiments.
@@ Appendix F Quantitative Analysis of Flaws in Chatbot Arena’s Evaluation Practice
We then examine the trend of human voter preferences when comparing two responses, one providing a toxic response and the other providing a non-toxic response (often by refusing to answer). We focus on human comparison pairs where one response is flagged as toxic and the other is not. Across all such comparison pairs, the proportion where the toxic response is preferred is 50.89%; when also counting “tie” cases to consider all cases where the non-toxic response is not preferred, the proportion rises to 71.98%. This issue is even more serious at the node “Facilitating inclusive, ethical, and strategic communication and engagement across diverse and sensitive contexts”; among comparison pairs for the node’s instructions, these two numbers rise significantly to 86.84% and 97.37%, respectively. These results confirm the observation that human voters tend to prefer toxic responses (that do not refuse to answer), diverging from the intended evaluation goals and values. They underscore the need for careful refinement of evaluation practices to ensure alignment with the desired principles.
@@ Appendix G Ablation Study: Alternative Approach to Tree Construction
Table 33: Accuracy (%) of different LMs on MATH and DS-1000 test sets. The initial LM is Llama 3.1 8B Instruct (Dubey et al., 2024) for MATH and DeepSeek-Coder-Base 6.7B (Guo et al., 2024) for DS-1000, respectively. See Section 5.3 for the experimental setup. Here, we compare EvalTree using the default capability tree construction pipeline with EvalTree using the capability tree built with the hierarchical clustering algorithm. Synthetic data (used to train the initial LM) are generated with the guidance of the weakness profiles produced by the two versions of EvalTree, respectively. The accuracy (of a trained LM) is reported as mean±stderr (“stderr” refers to standard error) across five random seeds.

### v2

@@ E.3.2 Analysis on Experimental Results
To better understand why TextDiff and QualEval are outperformed, we show the Precision and Recall curves in Figure 11 and 12. These curves show that both methods suffer from poor Precision, indicating that the weaknesses they identify cannot precisely pinpoint where the LM fails. We present the identified weaknesses from TextDiff, QualEval, and EvalTree when the weakness profile size is 10 in Table 28, along with their corresponding Precision, Recall, and F1; they are based on applying the three methods to the synthetic evaluation result generated for the MATH benchmark, with the probability hyperparameters set to $p=0.7$ and $d=0.2$. We observe that EvalTree achieves significantly higher Precision compared to the other two methods, while maintaining a quite high Recall, indicating that EvalTree can more precisely pinpoint specific areas where the LM underperforms and thus better satisfy desideratum 1. For example, EvalTree identified the weakness “Analyzing and applying relationships among polynomial expressions and their roots using Vieta’s formulas,” which closely aligns with the ground-truth weakness “Solving polynomial equations by analyzing relationships through Vieta’s formulas;” in contrast, TextDiff and QualEval identified two much coarser-grained weaknesses, “Handling problems involving the properties of polynomials and their roots” and “Solving linear, polynomial, and quadratic equations, including factoring and roots” respectively, failing to capture the critical aspect of Vieta’s formulas.
@@ E.3.3 Computing F1 on a Separate Set
In this subsection, we present the results of Section 5.2 using a separate test set (distinct from the profiling set) for computing $A$ in the formulas provided in Appendix E.3.1.
@@
Figure 13: F1 score curves of TextDiff, QualEval, and EvalTree, with the weakness profile size varying from 1 to 20. Precision, Recall, and thus F1 (more specifically, $A$ in the formulas provided in Appendix E.3.1) are computed on a separate test set, distinct from the profiling set used to generate the synthetic evaluation results. A horizontal line indicates each method’s highest score. $d$ is a hyperparameter to control the sampling probability.
@@ E.4 Experimental Details of Extrinsic Assessment
We use OpenAI’s gpt-4o-mini-2024-07-18 (OpenAI, 2024a) in our experiments to generate (synthetic) data inputs; the input generation prompts for MATH (Hendrycks et al., 2021b) and DS-1000 (Lai et al., 2023) are provided in Table 29 and Table 30, respectively; we set the max new tokens and temperature to 4096 and 1.0 (for generation diversity), respectively. We also use gpt-4o-mini-2024-07-18 to generate outputs for each collected input; the output generation prompts for MATH and DS-1000 are provided in Table 31 and Table 32, respectively; we set the max new tokens and temperature to 4096 and 0.0, respectively.
@@
For the EvalTree-guided data collection strategy, we set the accuracy threshold $\tau$ to 0.4 in the node extraction algorithm described in Section 3.2. This resulted in 9 identified weaknesses for MATH and 5 for DS-1000; the same number of weaknesses was identified when using the TextDiff-guided strategy and the QualEval-guided strategy, ensuring that all weakness-guided data collection strategies use weakness profiles of the same size. When sampling five in-context examples for input generation given an identified weakness in a weakness-guided data collection strategy, the examples are sampled from the associated instances (in the profiling set) of the identified weakness in the TextDiff-guided strategy, from the instances assigned to the identified weakness in the QualEval-guided strategy, and from the instances linked to the corresponding node in the EvalTree-guided strategy.
We provide an example of synthetic data inputs generated for Llama 3.1 8B Instruct on MATH. One EvalTree-identified weakness is “Analyzing and optimizing geometric relationships using trigonometric principles and the Triangle Inequality.” A synthetic data input generated under the guidance of this weakness is “In triangle $ABC$, the lengths of sides $AB$ and $AC$ are 15 cm and 20 cm, respectively. If angle $A$ measures $60^{\circ}$, what is the length of side $BC$ rounded to the nearest whole number?” In contrast, a synthetic data input guided by the generic capability is “A trader bought a certain number of apples for \$0.75 each and then sold them for \$1.00 each. If he had a total profit of \$15 after selling all the apples, how many apples did he sell?” This example highlights that EvalTree provides targeted guidance for data collection.
For each data collection strategy, we collect 128 instance inputs for training. We finetune the models using LoRA (Hu et al., 2022), with a rank of 256, an alpha of 512, and a dropout rate of 0.1. The batch size is fixed at 8, and the maximum sequence length is set to 1024 tokens. Training is conducted using BF16 precision. The optimizer is configured with a learning rate of 1E-4, a cosine learning rate scheduler, a warmup ratio of 0.1, and no weight decay. The models are trained for 3 and 2 epochs in the experiments on MATH and DS-1000, respectively. These configurations are applied consistently across all experiments.
@@ Appendix F Quantitative Analysis of Flaws in Chatbot Arena’s Evaluation Practice
We then examine the trend of human voter preferences when comparing two responses, one providing a toxic response and the other providing a non-toxic response (often by refusing to answer). We focus on human comparison pairs where one response is flagged as toxic and the other is not. Across all such comparison pairs, the proportion where the toxic response is preferred is 50.89%; when also counting “tie” cases to consider all cases where the non-toxic response is not preferred, the proportion rises to 71.98%. This issue is even more serious at the node “Facilitating inclusive, ethical, and strategic communication and engagement across diverse and sensitive contexts”; among comparison pairs for the node’s instructions, these two numbers rise significantly to 86.84% and 97.37%, respectively. These results confirm the observation that human voters tend to prefer toxic responses (that do not refuse to answer), diverging from the intended values. They underscore the need for careful refinement of evaluation practices to ensure alignment with the desired principles.
@@ Appendix G Ablation Study: Alternative Approach to Tree Construction
Table 33: Accuracy (%) of different LMs on MATH and DS-1000 test sets. The initial LM is Llama 3.1 8B Instruct (Dubey et al., 2024) for MATH and DeepSeek-Coder-Base 6.7B (Guo et al., 2024) for DS-1000, respectively. See Section 5.3 for the experimental setup. We compare EvalTree using the default capability tree construction pipeline with EvalTree using the capability tree built with the hierarchical clustering algorithm here. Synthetic data (used to train the initial LM) are generated under the guidance of the weakness profiles produced by the two versions of EvalTree, respectively. The accuracy (of a trained LM) is reported as mean±stderr (“stderr” refers to standard error) across five random seeds.